An administrator of a popular web site is told that a new server can handle 11,000 "hits" (users accessing the site) per second. The web site currently experiences a peak demand of about 85,000 hits per second; but every month, the peak demand increases by 3500 hits per second. Use a proportion equation to determine how many new servers the administrator should buy to address expected traffic for the next 18 months.
Answer by stanbon(75887)

An administrator of a popular web site is told that a new server can handle 11,000 "hits" (users accessing the site) per second. The web site currently experiences a peak demand of about 85,000 hits per second; but every month, the peak demand increases by 3500 hits per second. Use a proportion equation to determine how many new servers the administrator should buy to address expected traffic for the next 18 months.
--------------
x/(18*3500) = 1/11000
x = 63000/11000
x = 5.72727272..
---------
He will need 6 new servers
==============
Cheers,
Stan H.